OpenStudy (anonymous):

What is the hang time for a football that is kicked at 22.0 m/s at an angle of 45 degrees above the horizontal? What is the formula?

OpenStudy (anonymous):

i cant wait for this :-0

OpenStudy (anonymous):

Only the vertical motion decides hang time (how long ball is in the air). Vertical component of initial velocity = 22*sin(45) = 15.55 m/s final velocity (at the highest point of the ball) = 0 m/s acceleration = -9.8 m/s^2 Now (final velocity-initial velocity)=aXt this gives us 0-15.55=(-9.8)Xt which gives us t=1.586 seconds Ball takes 1.586 second to go up, and same time to come down. So total time in air is 1.586*2=3.173 seconds

OpenStudy (anonymous):

Do you always calculate the hang time with sin? Because all of the ones that I have seen have used sin

OpenStudy (anonymous):

Yes, you need sine of the angle with the horizontal to find the vertical component. hang time is decided by the vertical component, and sine of angle with the horizontal will give you that

OpenStudy (anonymous):

That is if the angle given to you is the angle with the horizontal ... otherwise you may need cosine or some other thing ... depends on what angle they give you as well.

OpenStudy (anonymous):

okay that makes sense. What kind of angle would need a cosine?

OpenStudy (anonymous):

if the give you the angle with the vertical instead of the horizontal

OpenStudy (anonymous):

Okay awesome. I am going to write that down in my notes. Thanks!

OpenStudy (anonymous):

Anytime